# BevSplat: Resolving Height Ambiguity via Feature-Based Gaussian Primitives for Weakly-Supervised Cross-View Localization

### v2

@@ Abstract
This paper addresses the problem of weakly supervised cross-view localization, where the goal is to estimate the pose of a ground camera relative to a satellite image with noisy ground truth annotations. A common approach to bridge the cross-view domain gap for pose estimation is Bird’s-Eye View (BEV) synthesis. However, existing methods struggle with height ambiguity due to the lack of depth information in ground images and satellite height maps. Previous solutions either assume a flat ground plane or rely on complex models, such as cross-view transformers. We propose BevSplat, a novel method that resolves height ambiguity by using feature-based Gaussian primitives. Each pixel in the ground image is represented by a 3D Gaussian with semantic and spatial features, which are synthesized into a BEV feature map for relative pose estimation. We validate our method on the widely used KITTI and VIGOR datasets, which include both pinhole and panoramic query images. Experimental results show that BevSplat significantly improves localization accuracy over prior approaches.
@@ 2.1 Cross-view Localization
Cross-view localization, the task of aligning ground-level images with satellite imagery, has become increasingly important in localization algorithms. Early approaches framed this as an image retrieval problem, where ground images were matched with satellite image slices of regions such as urban areas. Metric learning methods were used to train feature representations, enabling similarity computation between query ground images and satellite slices, facilitating localization [8, 9, 10, 11, 12]. With the advent of complex models like transformers, cross-view localization based on image retrieval has shown improved performance on slice databases, though practical application remains challenging [13, 14].
Recognizing these limitations, [15] introduced the one-to-many cross-view localization task. Building on this, recent works [16, 17, 18, 3, 19, 5, 20, 21, 22] advanced pixel-level localization methods. However, these approaches often assume precise pose information in the training data, which is typically derived from GPS signals and prone to inaccuracies in real-world deployment. To overcome this, [1, 2] proposed weakly supervised settings with noisy pose annotations. Note that [2] assumes the availability of GT labels in the source domain training dataset and cross-view image pairs in the target domain for training. In contrast, [1] addresses the more challenging scenario where GT labels are unavailable in the source domain training dataset, and no cross-view image pairs accessible in the target domain. We tackle the same task as [1].
@@ 2.2 Bird’s-Eye View Synthesis
BEV synthesis, which generates bird’s-eye view images from ground-level perspectives, has been widely applied to cross-view localization. While LiDAR and Radar sensors offer high accuracy for localization tasks [23, 24, 25, 26], their high cost limits their use. For camera-only systems, multi-camera setups are commonly employed [27, 28, 29, 30], primarily focusing on tasks like segmentation and recognition. In localization, methods like Inverse Perspective Mapping (IMP) assume a flat ground plane for BEV synthesis [1, 6], which can be overly simplistic for complex environments. Transformer-based models address these challenges but struggle with weak supervision and noisy pose annotations [3, 4, 5]. While methods such as [31, 32] also employ feature Gaussians for the image-to-BEV transformation, they typically benefit from rich depth and semantic information afforded by multi-sensor setups. While effective in some contexts, they face limitations in resource-constrained, real-world scenarios. In stark contrast, our approach is constrained to rely exclusively on weakly supervised signals derived purely from images, presenting a considerably more challenging task.
@@ 2.3 Sparse-View 3D Reconstruction
In our method, we adopt algorithms similar to 3D reconstruction to represent ground scenes. Sparse-view 3D reconstruction has been a major focus of the community. Nerf-based approaches [33] and their adaptations [34] have shown the potential for single-view 3D reconstruction, though their application is limited by small-scale scenes and high computational cost. Recent works using diffusion models [35] and 3D Gaussian representations [7][36, 37, 38], as well as transformer- and Gaussian-based models[39, 40], have achieved sparse-view 3D reconstruction on a larger scale, but the complexity of these models still restricts their use due to computational demands. Approaches like [41, 42, 43] leverage pre-trained models to directly generate Gaussian primitives, avoiding the limitations of complex models while enabling scene reconstruction from sparse views. We apply such methods to single-view reconstruction, achieving high-accuracy cross-view localization.
@@ 3 Method
In this paper, we address cross-view localization by aligning ground-level and satellite images under weak supervision, where initial ground camera locations are only approximate. Our objective is to accurately estimate camera pose from these noisy priors by leveraging Gaussian primitives, which effectively manage height ambiguity and enable efficient generation of Bird’s-Eye View (BEV) feature maps.
@@ 3.2.2 Pose Estimation via Confidence-Guided Similarity Learning
[…]
@@ 4 Experiments
KITTI dataset. The KITTI dataset [49] consists of ground-level images captured by a forward-facing pinhole camera with a restricted field of view, complemented by aerial images [50], where each aerial patch covers a ground area of approximately $100\times 100\mathrm{m}^{2}$. The dataset includes a training set and two test sets: Same-Area contains images from the same region as the training set, while Cross-Area consists of images from a different region. The location search range of ground images is approximately $56\times 56\mathrm{m}^{2}$, with an orientation noise of ±$\pm 10^{\circ}$.
VIGOR dataset. The VIGOR dataset [15] includes geo-tagged ground panoramas and satellite images from four US cities: Chicago, New York, San Francisco, and Seattle. Each satellite patch spans $70\times 70\mathrm{m}^{2}$ and is labeled positive if the ground camera is within its central $1/4$ region; otherwise, it is semi-positive. The dataset has Same-Area and Cross-Area splits: Same-Area uses training and testing data from the same region, while Cross-Area splits training and testing between two separate city groups. We use only positive satellite images for all experiments, following [1].
@@
Implementation Details. For 3D point cloud generation from ground images, we employ specific depth estimation models: DepthAnythingV2 [51] for the pinhole camera images of the KITTI dataset [49], and UniK3D [52] for the panoramic images of the VIGOR dataset [15]. Our feature extractor for both ground and satellite images is a DINOv2 backbone [47], initialized with FiT weights [46], which is subsequently fine-tuned using an attached DPT module [48]. This extractor yields satellite feature maps with dimensions $(C,H,W)=(32,128,128)$. For ground images, initial feature maps of $(32,64,256)$ are extracted; these are then projected into BEV using our feature Gaussians projection, resulting in final ground BEV features also of dimensions $(32,128,128)$. Our model is trained using the AdamW optimizer [53] (weight decay $10^{-3}$) and a OneCycleLR scheduler [54] with a cosine annealing strategy, where the learning rate peaks at $6.25\times 10^{-5}$. We use a batch size of 8 on a single 4090 GPU, training for 6 epochs on KITTI and 10 epochs on VIGOR.
@@ 4.1 Comparison with State-of-the-Art Methods
We compare our method with the latest state-of-the-art (SOTA) approaches, including supervised methods such as Boosting [4], CCVPE [55], HC-Net [6], and DenseFlow [18], all of which rely on ground-truth camera poses for supervision. We also compare with G2Sweakly [1], which uses only a satellite image and a corresponding ground image as input, similar to our setup.
@@
Our feature-based Gaussian splatting for BEV synthesis effectively addresses these issues, leading to a notable enhancement in localization accuracy. Fig. 1 and Fig. 4 visualize the difference between the IPM projection and our proposed BEV synthesis method, clearly demonstrating that our projection technique resolves challenges such as occlusions caused by tall objects (e.g., buildings, trees, vehicles) and geometric distortions from curved roads. Furthermore, in cross-area evaluations, our method even surpasses supervised approaches (Boosting [4], CCVPE [55], HC-Net [6], and DenseFlow [18]) in terms of mean and median errors, showcasing the strong generalization ability of our approach and highlighting the potential of weakly supervised methods.
@@
Algorithms $\lambda_{1}$ Test Area localization Lateral Longitudinal Azimuth mean(m)$\downarrow$ median(m)$\downarrow$ d=1$\uparrow$ d=3$\uparrow$ d=1$\uparrow$ d=3$\uparrow$ $\theta=1^{\circ}$$\uparrow$ $\theta=3^{\circ}$$\uparrow$ mean($\downarrow$ median($\downarrow$ Boosting [4]* - Same-Area - - 76.44 96.34 23.54 50.57 99.10 100.00 - - CCVPE [55]* - 1.22 0.62 97.35 98.65 77.13 96.08 77.39 99.47 0.67 0.54 HC-Net [6]* - 0.80 0.50 99.01 - 92.20 - 91.35 99.84 0.45 0.33 DenseFlow [18]* - 1.48 0.47 95.47 - 87.89 - 89.40 - 0.49 0.30 G2SWeakly [1] 0 12.03 8.10 59.58 85.74 11.37 31.94 99.99 100.00 0.33 0.28 Ours 0 5.82 2.85 60.04 91.54 24.06 56.82 99.99 100.00 0.33 0.28 G2SWeakly [1] 1 6.81 3.39 66.07 94.22 16.51 49.96 99.99 100.00 0.33 0.28 Ours 1 2.87 2.06 60.28 94.24 35.62 76.57 99.99 100.00 0.33 0.28 Boosting [4]* - Cross-Area - - 57.72 86.77 14.15 34.59 98.98 100.00 - - CCVPE [55]* - 9.16 3.33 44.06 81.72 23.08 52.85 57.72 92.34 1.55 0.84 HC-Net [6]* - 8.47 4.57 75.00 - 58.93 - 33.58 83.78 3.22 1.63 DenseFlow [18]* - 7.97 3.52 54.19 - 23.10 - 43.44 - 2.17 1.21 G2SWeakly [1] 0 13.87 10.24 62.73 86.53 9.98 29.67 99.99 100.00 0.33 0.28 Ours 0 7.05 3.22 63.06 92.62 23.08 51.61 99.99 100.00 0.33 0.28 G2SWeakly [1] 1 12.15 7.16 64.74 86.18 11.81 34.77 99.99 100.00 0.33 0.28 Ours 1 6.20 2.51 60.01 95.17 27.41 60.45 99.99 100.00 0.33 0.28
@@
Method $\lambda_{1}$ Same-Area Cross-Area Aligned-orientation Unknown-orientation Aligned-orientation Unknown-orientation mean(m)$\downarrow$ median(m)$\downarrow$ mean(m)$\downarrow$ median(m)$\downarrow$ mean(m)$\downarrow$ median(m)$\downarrow$ mean(m)$\downarrow$ median(m)$\downarrow$ Boosting [4]* - 4.12 1.34 - - 5.16 1.40 - - CCVPE [55]* - 3.60 1.36 3.74 1.42 4.97 1.68 5.41 1.89 HC-Net [6]* - 2.65 1.17 - - 3.35 1.59 - - DenseFlow [18]* - 3.03 0.97 4.97 1.90 5.01 2.42 7.67 3.67 G2SWeakly [1] 0 5.22 1.97 5.33 2.09 5.37 1.93 5.37 1.93 Ours 0 2.96 1.62 2.98 1.74 2.90 1.65 2.93 1.73 G2SWeakly [1] 1 4.19 1.68 4.18 1.66 4.70 1.68 4.52 1.65 Ours 1 2.57 1.40 2.61 1.41 2.63 1.38 2.67 1.46
@@
Computation comparison. All evaluations of GPU memory usage were performed on an NVIDIA RTX 4090. Our model, which features a DINOv2 [47] backbone fine-tuned with our lightweight DPT network [48] (composed of a few CNN layers), requires considerably less memory during the training phase compared to G2SWeakly [1]. On the KITTI [49] dataset (batch size 8), our training phase uses only 9.2,GB of GPU memory, substantially less than the 22.7,GB required by G2SWeakly [1]. During inference, our model consumes 7.7,GB of GPU memory, while G2SWeakly [1] requires 7.2,GB. This increase 0.5,GB for our method is attributable to the larger DINOv2 backbone [47], representing a trade-off for its enhanced feature representation capabilities.
Table 3: BEV synthesis comparison on KITTI dataset.
@@
Direct Projection 1 4.43 3.29 7.94 4.64
@@
Table 4: Ablation study on backbone module on KITTI dataset.
@@
VGG [56] 0 8.77 4.53 9.91 5.80
DINOv1 [57] 0 7.68 4.01 9.16 4.67
DINOv2 [47] 0 7.04 3.45 8.37 4.21
DINOv2(FIT) [46] 0 5.82 2.85 7.05 3.22
VGG [56] 1 6.49 2.72 8.02 4.29
DINOv1 [57] 1 4.77 2.98 7.12 3.37
DINOv2 [47] 1 4.21 2.73 7.01 3.18
DINOv2(FIT) [46] 1 2.87 2.06 6.20 2.51
4.2 Ablation Study
@@
The IPM projection method. This technique assumes all pixels in the ground-level image correspond to a flat plane at a real-world height of 0 meters. Consequently, while IPM can accurately represent flat road surfaces in BEV, it introduces significant distortions for any objects with non-zero elevations. These objects are typically stretched along the line of sight in the BEV. For instance, in the first example depicted in Fig. 4, vehicles’ features appear elongated into regions not corresponding to their actual ground footprint. Similarly, in the second example (Fig. 4), buildings’ features are distorted and erroneously projected. A further limitation is that IPM typically projects only the lower portion of the ground image to BEV, discarding valuable information from the upper half. In contrast, our BevSplat method is designed to leverage geometric information from the entire image more effectively.
@@
Foundation model backbone. Prioritizing robust outdoor generalization and effective 3D-relevant feature extraction for our foundation model, we selected DINOv2 [47] fine-tuned with the FiT method [46]. This model, which we term DINOv2(FiT), utilizes its dinov2_base_fine pre-trained weights renowned for these capabilities. To validate this choice and compare its efficacy against alternatives, our ablation study also evaluated VGG [56], DINOv1 [57], and the original DINOv2. All DINO-based backbones in this study (DINOv1, DINOv2, and DINOv2(FiT)) were subsequently further fine-tuned by us using a DPT-like module [48]. The ablation results (Table 4) validated our selection, as DINOv2(FiT), after our DPT-like fine-tuning, demonstrated superior performance among the evaluated backbones.

### v3

@@ Abstract
This paper addresses the problem of weakly supervised cross-view localization, where the goal is to estimate the pose of a ground camera relative to a satellite image with noisy ground truth annotations. A common approach to bridge the cross-view domain gap for pose estimation is Bird’s-Eye View (BEV) synthesis. However, existing methods struggle with height ambiguity due to the lack of depth information in ground images and satellite height maps. Because a single 2D pixel could represent points at various depths and heights, its true 3D position is ambiguous. Previous solutions either assume a flat ground plane or rely on complex models, such as cross-view transformers. We propose BevSplat, a novel method that resolves height ambiguity by using feature-based Gaussian primitives. Each pixel in the ground image is represented by a 3D Gaussian with semantic and spatial features, which are synthesized into a BEV feature map for relative pose estimation. We validate our method on the widely used KITTI and VIGOR datasets, which include both pinhole and panoramic query images. Experimental results show that BevSplat significantly improves localization accuracy over prior approaches.
@@ 2.1 Cross-view Localization
Cross-view localization, which aligns ground-level images with satellite imagery, has evolved from image retrieval to fine-grained pose estimation. Early approaches framed this as an image retrieval task, using metric learning to match ground queries to satellite image slices [8, 9, 10, 11, 12]. While modern transformers have improved retrieval performance, practical application remains challenging [13, 14]. Many recent methods adopt a coarse-to-fine pipeline. This typically involves a coarse retrieval step [15] followed by fine-grained (pixel-level) localization to identify the precise camera pose [16, 17, 18, 3, 19, 5, 20, 21, 22]. A key limitation of these methods is their reliance on precise, GPS-based training data, which is often prone to inaccuracies. To overcome this, weakly supervised settings have been proposed to learn from noisy pose annotations [1, 2]. These weakly supervised settings differ in their assumptions. [2] assumes the availability of GT labels in a source domain and access to cross-view pairs in the target domain. In contrast, [1] addresses a more challenging scenario where source domain GT labels are unavailable and no target domain pairs are accessible. In this work, we tackle the same task setting as [1].
@@ 2.2 Bird’s-Eye View Synthesis
BEV synthesis, which generates bird’s-eye view images from ground-level perspectives, has been widely applied to cross-view localization. While LiDAR and Radar sensors offer high accuracy for localization tasks [23, 24, 25, 26], their high cost limits their use. For camera-only systems, multi-camera setups are commonly employed [27, 28, 29, 30], primarily focusing on tasks like segmentation and recognition. In localization, methods like Inverse Perspective Mapping (IMP) assume a flat ground plane for BEV synthesis [1, 6, 20, 18], which can be overly simplistic for complex environments. Transformer-based models address these challenges but struggle with weak supervision and noisy pose annotations [3, 4, 5]. While methods such as [31, 32] also employ feature Gaussians for the image-to-BEV transformation, they typically benefit from rich depth and semantic information afforded by multi-sensor setups. While effective in some contexts, they face limitations in resource-constrained, real-world scenarios. In stark contrast, our approach is constrained to rely exclusively on weakly supervised signals derived purely from images, presenting a considerably more challenging task.
@@ 2.3 Sparse-View 3D Reconstruction
In our method, we adopt algorithms similar to 3D reconstruction to represent ground scenes. Sparse-view 3D reconstruction has been a major focus of the community. Nerf-based approaches [33] and their adaptations [34] have shown the potential for single-view 3D reconstruction, though their application is limited by small-scale scenes and high computational cost. Recent works using diffusion models [35, 36, 37] and 3D Gaussian representations [7, 38, 39, 40], as well as transformer- and Gaussian-based models[41, 42], have achieved sparse-view 3D reconstruction on a larger scale, but the complexity of these models still restricts their use due to computational demands. Approaches like [43, 44, 45] leverage pre-trained models to directly generate Gaussian primitives, avoiding the limitations of complex models while enabling scene reconstruction from sparse views. We apply such methods to single-view reconstruction, achieving high-accuracy cross-view localization.
@@ 3 Method
In this paper, we address cross-view localization by aligning ground-level and satellite images under weak supervision, where initial ground camera locations are only approximate. Our objective is to accurately estimate camera pose from these noisy priors by leveraging Gaussian primitives, which effectively manage height ambiguity and enable efficient generation of Bird’s-Eye View (BEV) feature maps. First, we employ an orientation prediction network analogous to that in G2SWeakly to align the orientations of the ground and satellite images. Subsequently, our BevSplat method lifts the ground image to 3D (Section 3.1) and projects the corresponding Feature Gaussians into the BEV perspective to render the ground-view BEV features (Section 3.2.1). Finally, these features are compared against the satellite features by computing a similarity score (Section 3.2.2).
@@ 3.2.2 Pose Estimation via Confidence-Guided Similarity Learning
Here, $\mathbf{F}_{s}$ denotes the satellite image features, while $\mathbf{F}_{BEV}$ and $\mathbf{C}_{BEV}$ are the BEV features and confidence map derived from the ground image, respectively. $\|\cdot\|$ signifies the $L_{2}$ norm.
@@ 4 Experiments
KITTI dataset. The KITTI dataset [51] consists of ground-level images captured by a forward-facing pinhole camera with a restricted field of view, complemented by aerial images [52], where each aerial patch covers a ground area of approximately $100\times 100\mathrm{m}^{2}$. The dataset includes a training set and two test sets(Same-Area and Cross-Area). For the Same-Area test set, the test query images are from the same geographical regions as the training set, but are not the same images. For the Cross-Area test set, the test images come from entirely new geographical regions that were not seen during training, testing the model’s ability to generalize. The location search range of ground images is approximately $56\times 56\mathrm{m}^{2}$, with an orientation noise of $\pm 10^{\circ}$.
VIGOR dataset. The VIGOR dataset [15] includes geo-tagged ground panoramas and satellite images from four US cities: Chicago, New York, San Francisco, and Seattle. Each satellite patch spans $70\times 70\mathrm{m}^{2}$ and is labeled positive if the ground camera is within its central $1/4$ region; otherwise, it is semi-positive. The dataset also has Same-Area and Cross-Area splits: Same-Area uses training and testing data from the same region, while Cross-Area splits training and testing between two separate city groups. We use only positive satellite images for all experiments, following [1].
@@
Implementation Details. For 3D point cloud generation from ground images, we employ specific depth estimation models: DepthAnythingV2 [53] for the pinhole camera images of the KITTI dataset [51], and UniK3D [54] for the panoramic images of the VIGOR dataset [15]. Our feature extractor for both ground and satellite images is a DINOv2 backbone [49], initialized with FiT weights [48], which is subsequently fine-tuned using an attached DPT module [50]. This extractor yields satellite feature maps with dimensions $(C,H,W)=(32,128,128)$. For ground images, initial feature maps of $(32,64,256)$ are extracted; these are then projected into BEV using our feature Gaussians projection, resulting in final ground BEV features also of dimensions $(32,128,128)$. Our model is trained using the AdamW optimizer [55] (weight decay $10^{-3}$) and a OneCycleLR scheduler [56] with a cosine annealing strategy, where the learning rate peaks at $6.25\times 10^{-5}$. We use a batch size of 8 on a single 4090 GPU, training for 8 epochs on KITTI and 14cf epochs on VIGOR.
@@ 4.1 Comparison with State-of-the-Art Methods
We compare our method with the latest state-of-the-art (SOTA) approaches, including supervised methods such as Boosting [4], VFA [20], CCVPE [57], HC-Net [6], and DenseFlow [18], all of which rely on ground-truth camera poses for supervision. We also compare with G2Sweakly [1], which uses only a satellite image and a corresponding ground image as input, similar to our setup.
@@
Our feature-based Gaussian splatting for BEV synthesis effectively addresses these issues, leading to a notable enhancement in localization accuracy. Fig. 1 and Fig. 4 visualize the difference between the IPM projection and our proposed BEV synthesis method, clearly demonstrating that our projection technique resolves challenges such as occlusions caused by tall objects (e.g., buildings, trees, vehicles) and geometric distortions from curved roads. Furthermore, in cross-area evaluations, our method even surpasses supervised approaches (Boosting [4], VFA [20], CCVPE [57], HC-Net [6], and DenseFlow [18]) in terms of mean and median errors, showcasing the strong generalization ability of our approach and highlighting the potential of weakly supervised methods.
It is worth noting that the experimental results for VFA [20] are taken from the PIDLoc [22] paper. This is because the authors of VFA have adopted a setting in their prior works [58, 59] that aligns ground-truth poses to the satellite image center, which risks overfitting by biasing predictions towards the center. Since the official code for VFA [20] is not available, we attempted to reproduce their results and found that we could only achieve the reported performance by using this same setting. However, when we use the standard setting adopted by our method and other comparable works [4, 57, 6, 18, 1], our reproduced results are consistent with the VFA results reported in PIDLoc [22]. Therefore, for a fair comparison, we use the VFA [20] results from PIDLoc [22].
@@
[…]
@@
[…]
@@
Computation comparison. All evaluations of GPU memory usage were performed on an NVIDIA RTX 4090. Our model, which features a DINOv2 [49] backbone fine-tuned with our lightweight DPT network [50] (composed of a few CNN layers), requires considerably less memory during the training phase compared to G2SWeakly [1]. On the KITTI [51] dataset (batch size 8), our training phase uses only 9.2 GB of GPU memory, substantially less than the 22.7 GB required by G2SWeakly [1]. During inference, our model consumes 7.7 GB of GPU memory, while G2SWeakly [1] requires 7.2 GB as shown in Table 5. This increase 0.5 GB for our method is attributable to the larger DINOv2 backbone [49], representing a trade-off for its enhanced feature representation capabilities.
Table 3: BEV synthesis comparison on the KITTI dataset.
@@
Lift-Splat-Shoot [60] 1 7.89 4.30 11.63 5.31
OrienterNet [5] 1 5.71 3.20 10.02 5.07
Direct Projection 1 7.59 4.25 8.93 5.81
@@
Table 4: Ablation study on backbone module on the KITTI dataset.
Methods $\lambda_{1}$ Same Area Cross Area
Direct Train 0 17.74 15.61 17.59 15.71
LoRA [61] 0 16.29 14.48 17.05 14.7
DPT 0 5.82 2.85 7.05 3.22
Direct Train 1 14.32 12.43 17.28 15.14
LoRA [61] 1 13.58 11.79 16.81 14.63
DPT 1 2.87 2.06 6.20 2.51
Table 5: Comparison of resource consumption.
Method Training Memory Inference Memory Inference Time
OrienterNet [5] $32.4$ $10.8$ $71$
LSS [60] $26.1$ $8.3$ $85$
G2SWeakly [1] $22.7$ $7.2$ $31$
Ours $9.2$ $7.7$ $44$
4.2 Ablation Study
Table 6: Ablation study on backbone module on the KITTI dataset.
@@ 4.2 Ablation Study
$downarrow Mean(m)\downarrow Median(m)\downarrow$
VGG 0 8.77 4.53 9.91 5.80
DINOv1 0 7.68 4.01 9.16 4.67
DINOv2 0 7.04 3.45 8.37 4.21
DINOv2(FIT) 0 5.82 2.85 7.05 3.22
VGG 1 6.49 2.72 8.02 4.29
DINOv1 1 4.77 2.98 7.12 3.37
DINOv2 1 4.21 2.73 7.01 3.18
DINOv2(FIT) 1 2.87 2.06 6.20 2.51
@@
The IPM projection method. This technique assumes all pixels in the ground-level image correspond to a flat plane at a real-world height of 0 meters. Consequently, while IPM can accurately represent flat road surfaces in BEV, it introduces significant distortions for any objects with non-zero elevations. These objects are typically stretched along the line of sight in the BEV. For instance, in the first row depicted in Fig. 4(b), vehicles’ features appear elongated into regions not corresponding to their actual ground footprint. Similarly, in the second row Fig. 4(b), buildings’ features are distorted and erroneously projected. A further limitation is that IPM typically projects only the lower portion of the ground image to BEV, discarding valuable information from the upper half. In contrast, our BevSplat method is designed to leverage geometric information from the entire image more effectively.
@@
Other depth-based re-sampling methods. We further compare our method with other BEV projection techniques that are also based on depth prediction, such as those used in Lift-Splat-Shoot [60] and OrienterNet [5] as detailed in Table 5, by adapting their projection modules into our framework. Although these approaches can generate a denser BEV and, like our method, leverage the full vertical information from the ground-view image—allowing them to perform well in same-area settings when guided by GPS labels—they fail to generalize to unseen environments. They tend to make erroneous guesses to fill in occluded regions, which explains their performance degradation in cross-area evaluations. Furthermore, they employ a simple weighted averaging for BEV projection, which is less accurate for handling vertical occlusions compared to BevSplat’s principled alpha blending. Finally, both methods utilize a complex $h\times w\times d$ depth representation to perform an attention-based sum over ground features. This implicit, high-dimensional process incurs substantial computational and memory overhead to produce a denser BEV as shown in Table 5.
@@
Foundation model backbone. To validate the effectiveness of fine-tuning a foundation model for extracting ground and satellite image features, we conducted ablation studies on the impact of different foundation models with their pre-trained weights, as well as the influence of various fine-tuning methods on the experimental results.
Impact of Different Foundation Models and Weights. Prioritizing robust outdoor generalization and effective 3D-relevant feature extraction for our foundation model, we selected DINOv2 [49] fine-tuned with the FiT method [48]. This model, which we term DINOv2(FiT), utilizes its dinov2_base_fine pre-trained weights renowned for these capabilities. To validate this choice and compare its efficacy against alternatives, our ablation study also evaluated VGG [62], DINOv1 [63], and the original DINOv2. All DINO-based backbones in this study (DINOv1, DINOv2, and DINOv2(FiT)) were subsequently further fine-tuned by us using a DPT-like module [50]. The ablation results (Table 6) validated our selection, as DINOv2(FiT), after our DPT-like fine-tuning, demonstrated superior performance among the evaluated backbones.
Impact of Different Fine-tuning Methods. Using DPT-like models is a common practice for 3D vision tasks; for example, VGGT [64] utilizes DPT [50] for point cloud reconstruction, depth estimation, and feature matching. Although we are the first to apply DPT-DINO for feature extraction in the specific sub-field of cross-view localization, our motivation is to similarly obtain features that are rich in 3D information. This is analogous to human navigation, where in addition to semantic information, an understanding of the real 3D scene is also crucial for localization. However, obtaining such 3D-aware features to bridge the significant ground-satellite domain gap is non-trivial. Simpler fine-tuning methods like direct end-to-end training or LoRA [61] fail, as they either lose crucial texture details or are not powerful enough to adapt the foundation model. We use a DPT-like module because its multi-scale feature fusion architecture is uniquely suited for this challenge. It successfully adapts the backbone by preserving both the low-level texture and high-level semantic information required for matching across these different domains as shown in Table 5.